A cricket ball of mass 100 gram is moving with a speed of 25 metre per second and brought to rest in 0.025 seconds find the average force apply and impulse

Grade:11

2 Answers

Arun
6 years ago
mass of the ball (m) =100g or 0.1kg
initial velocity (u) =25 m/s
time (t) = 0.025sec.
final velocity (v) = 0
  • Initial momentum = mu
                              = 0.1kg*25m/s​
                               = 2.5 kgm/s
  • Final momentum = mv (which is equal to 0, since v is zero)
  • Change in momentum = mv-mu
                                               = 0 - 2.5 = -2.5 kgm/s
  • Average force = Change in momentum / time
                                  = -2.5 kgms-1 /0.025s
                                  = -100 kgm/s2
                                 = -100 N

The negative sign actually shows that the force was applied oppasite to the direction of the motion of the ball. So, the average force  applied by the player will be 100 N.
anjace
6 years ago
impulse=f*t
f=ma
acceleration can be calculated by  1st lw of motion
a=-1000
avg force=impulse/time 
which is equal to -100 n